OpenStudy (anonymous):

Using a compass, Jane draws a circle on her paper with a radius of 2.5 inches. What is the area of the circle? A) 15.7 in2 B) 19.625 in2 C) 24.649 in2 D) 61.6225 in2

OpenStudy (anonymous):

\(A = \pi r^2\) Plug in what we know: \(A = (3.14)(2.5)^2\) Multiply 3.14 * 2.5 * 2.5

OpenStudy (anonymous):

(3.14)(2.5)(2.5) =19.625

OpenStudy (anonymous):

Yep, that's our answer.

OpenStudy (anonymous):

@iGreen. You are amazing... o-o

OpenStudy (anonymous):

xD Thanks